how do I evaluate -3/4^2 and write my answer as a fraction?

Found 4 solutions by MathLover1, MathTherapy, ikleyn, Alan3354:
Answer by MathLover1(20849)
Answer by MathTherapy(10552)


how do I evaluate -3/4^2 and write my answer as a fraction?
 1st thing: Use PARENTHESES! 


Answer by ikleyn(52781)

.
When you post the formulas to this forum

AND IF YOU WANT THE TUTORS UNDERSTAND YOUR FORMULAS, USE PARENTHESES.


It is a MANDATORY REQUIREMENT in this forum.




Answer by Alan3354(69443)

how do I evaluate -3/4^2 and write my answer as a fraction?
=========
(-3/4)^2 = 9/16
---
-(3/4)^2 = -9/16
---
-3/(4^2) = -3/16
---
You pick.